# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v2

@@ 3 Method
The task of diagnosing brain disorders primarily involves inferring specific properties (represented as class labels) from functional magnetic resonance imaging (fMRI) data in the form of graphs. Given a labeled dataset $\mathcal{D}=\{(\mathcal{G},\mathcal{Y})\}=\{(G_{i},y_{i})\}_{i=1}^{N}$, where each graph $G_{i}\in\mathcal{G}$ corresponds to a label $y_{i}\in\mathcal{Y}$, the goal is to extract features to train a mapping function $f_{\theta}:\mathcal{G}\rightarrow\mathcal{Y}$ that generalizes to unobserved graphs. For example, in studies of Alzheimer’s disease using brain networks, the label space comprises three diagnostic categories: NC (Normal Control), MCI (Mild Cognitive Impairment) and AD (Alzheimer’s Disease). The objective is to extract both lower-order and higher-order topological features from annotated brain network data in order to learn a robust classifier $f_{\theta}$, ensuring its effectiveness on novel, unseen brain connectomes.
@@ 3.1 Signed Higher-Order Interaction Representation Strategy
Most traditional network neuroscience approaches construct functional connectivity through pairwise Pearson correlations of fMRI time series (Fornito et al., 2016; Lindquist, 2008). Recent edge-centric frameworks account for interactions of edges by computing instantaneous co-fluctuations through element-wise products of z-scored signals $\mathbf{x}_{i}\odot\mathbf{x}_{j}$ (Faskowitz et al., 2020). Reference (Santoro et al., 2023, 2024) extended this approach to characterize $k$-order interactions when $k+1$ nodes simultaneously deviate from their temporal mean activation at time $t$. Although these techniques based on Pearson correlation have provided insights into the brain’s dynamic higher-order interactions, the methods have generally been limited by the lack of adequate temporal resolution (Shine et al., 2015). Now we give a new way to calculate co-fluctuations between several ROIs.
@@ 3.1.3 Construct two signed weighted simplicial complexes
It is important to acknowledge that in a weighted simplicial complex, denoted by $\mathcal{K}^{t}$, there exist two categories of signs for a simplex: concordant signs of positive and discordant signs of negative. The present study focuses exclusively on concordant signs of positive, eschewing the examination of discordant signs of negative. This is due to the potential for such signs to represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). The concordant interactions were subsequently categorised into two distinct groups: those that were positively synergistic and those that were negatively synergistic. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer valuable insights into the complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. As demonstrated in Fig. 2, it exhibits six distinct signed quadruplet higher-order interactions. For each temporal instance $t$, the weighted simplicial complexes are further extracted to construct two signed weighted simplicial complexes $\mathcal{K}^{t}_{positive}$ and $\mathcal{K}^{t}_{negative}$, namely the positive and negative complexes, respectively.
@@ 3.2.1 Filtrate the Weighted Simplicial Complexes
At each time point $t$, five distinct weighted networks are constructed: one derived from edge structures representing lower-order interactions, two from signed good tetrahedral structures signifying higher-order interactions, and another two from signed 2D holes illustrating the data’s intrinsic geometric shape from a higher-dimensional perspective. By averaging across the entire time period, the resulting networks exhibit enhanced stability and mitigate the impact of fMRI noise.
@@ 3.3.4 Attention-Guided Feature fusion Mechanism for Classification
For the embedding at the graph level ${F}_{i}$ of each channel obtained by flattening ${Z}^{G}_{i}$ ($i=1,2,3$), we igniting an attention-guided feature fusion mechanism to emphasize discriminative features while suppressing redundant information. The formula can be described as:
@@ 4.1.3 Implementation Details
    In HOI-Brain, we employ a Julia-based implementation of the Ripserer and PersistenceDiagrams libraries (Bauer, 2021) to compute persistent homology and homological scaffolds, extracting higher-order topological features.
@@ 4.2 Model Comparison
Table,2, 3, 4, 5 demonstrate the comparison results on ADNI, TaoWu, PPMI, and ABIDE datasets. The mean and standard deviation of ten-fold cross-validation were listed, with the best results highlighted in bold and the second-best results in underlining. Compared with traditional ML methods, general-purposed GNN-based models for brain network do not show a significant advantage and attain similar performance in most cases. This finding is consistent with the results reported in existing papers. This is because, under the same input conditions, neural network models typically have a larger number of parameters than conventional machine learning methods when utilizing edge features for better interpretability, making them more susceptible to overfitting. Transformer-based models predominantly outperform GNN-based models across all four datasets in most scenarios, highlighting their superior capability to capture the global structure of brain connectomes. HGNN-based models generally do not demonstrate superiority over other models, with notable performance only observed on the PPMI dataset. This phenomenon can be attributed to two main reasons: first, such models introduce excessive parameters, posing a risk of overfitting; second, their k-hop neighbor-based modeling approach fails to effectively capture the co-fluctuation characteristics among multiple brain regions, thereby inadequately representing higher-order interactions. In PH-based models, such as ATPGCN and Brain-HORs, which introduce higher-order loops or connected component features extracted via persistent homology for brain disease diagnosis, demonstrate significant superiority compared to both graph neural network (GNN)-based models and traditional machine learning models that rely solely on edge features. This highlights the necessity of incorporating higher-order topological features in brain disease diagnosis methods and provides a theoretical foundation for exploring brain diseases from higher-dimensional and more comprehensive perspectives.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNN-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating higher-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network Transformer, can effectively and holistically integrate complementary information from heterogeneous higher-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNN-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of higher-order interactions through the Multiplication of Temporal Derivatives and the extraction of topological features using persistent homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and voids (2D holes) formed by triangular faces and quadruplet.
@@
These experimental results demonstrate the effectiveness of our framework. To the best of our knowledge, our HOI-brain is the first method that distinguishes between positively and negatively synergistic higher-order interactions, and the first attempt to utilize quadruplet-level interaction signatures and two-dimensional void descriptors extracted via Persistent Homology to enhance diagnostic efficacy.
@@ 4.3.1 Ablation study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and voids (2D holes) formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extract lower-order edge features, violating triangles, one-dimensional loops. In addition, we extract unsigned good quadruplets and unsigned two-dimensional voids based our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Table 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be due to the fact that a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true higher-order interactions. Additionally, distinguishing between positively and negatively synergistic higher-order interactions to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may due to the fact that it can offer detailed insights into the communication within the brain.
@@
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network Transformer. The different variants are: not adding signed higher-order features decoupling mechanism (wo-signed), not adding orthonormal clustering readout of multiple channels, and not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Fig.4, we can see that adding signed higher-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain wins all the cases over the variant (wo-signed) or (wo-fusion). This result is consistent with our assumption that, by introducing the attention mechanism, our model can adaptively fuse positive and negative information as well as information from different channels, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both low-level and high-level feature patterns, which often reside within distinct functional modular structures.
@@ 4.4.1 In-depth Analysis of Attention Mechanism
Next, we also conducted a visualization analysis of the attention scores $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$ derived from the signed higher-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Fig. 7 shows that across both datasets, the importance of negative synergistic quadruplets is generally greater than that of positive synergistic information in most cases, while the importance of positive synergistic voids exceeds that of negative synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
In order to conduct a more in-depth investigation into the aforementioned phenomena in the attention mechanism, inter-group comparisons were conducted of the signed higher-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors. These were captured by the model in both the patient and healthy control groups. The ADNI dataset was used as an exemplar. As demonstrated in Fig. 8, which presents a comparison of group differences in higher-order topological features, it was observed that the group differences in the quadruplet-level interaction signatures were found to be highly significant, with all statistical significance levels at $p<0.001$ compared to void descriptors. This finding is consistent with the results of our model visualisation. This phenomenon may be attributed to the disruption of higher-order interactions within the brain during the progression of the disease, which in turn exerts an influence on the high-dimensional hole structures present within the brain. Furthermore, as the disease progresses, the number of positive quadruplets exhibits a gradual decrease, while the number of negative quadruplets initially decreases and subsequently increases. This finding suggests that during the transition from CN (cognitively normal) to MCI (mild cognitive impairment), there is a decline in synergistic interactions in the brain, whether positive or negative. Conversely, as the condition progresses from MCI to AD (Alzheimer’s disease), brain regions have been observed to exhibit negatively synergistic interactions, concurrently reducing their excitability to preserve the functionality of diverse neural modules. With regard to the number of voids, although no clear trend was observed in the progression of the disease, it was found that positive voids exhibited more significant intergroup differences compared to negative voids. This finding lends further support to the hypothesis that they exhibit higher levels of attention scores. Consequently, CN appears to utilise positive information to construct the whole-brain topological structure, whereas AD relies more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@ 4.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Fig. 9, 10. The attention scores are averaged across all subjects in the ADNI and ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both low-order and higher-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrices on the ADNI and ABIDE datasets in Fig. 11. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and higher-order patterns, which further demonstrates that varying degrees of modularity in both low-order and higher-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.

### v3

@@ 3 Method
The task of diagnosing brain disorders primarily involves inferring specific properties (represented as class labels) from fMRI data in the form of graphs. Given a labeled dataset $\mathcal{D}=\{(\mathcal{G},\mathcal{Y})\}=\{(G_{i},y_{i})\}_{i=1}^{N}$, where each graph $G_{i}\in\mathcal{G}$ corresponds to a label $y_{i}\in\mathcal{Y}$, the goal is to extract features to train a mapping function $f_{\theta}:\mathcal{G}\rightarrow\mathcal{Y}$ that generalizes to unobserved graphs. For example, in studies of Alzheimer’s disease using brain networks, the label space comprises three diagnostic categories: NC (Normal Control), MCI (Mild Cognitive Impairment) and AD (Alzheimer’s Disease). The objective is to extract both lower-order and higher-order topological features from annotated brain network data in order to learn a robust classifier $f_{\theta}$, ensuring its effectiveness on novel, unseen brain connectomes.
@@ 3.1 Signed Higher-Order Interaction Representation Strategy
Most traditional network neuroscience approaches construct functional connectivity through pairwise Pearson correlations of fMRI time series (Fornito et al., 2016; Lindquist, 2008). Recent edge-centric frameworks account for interactions of edges by computing instantaneous co-fluctuations through element-wise products of z-scored signals $\mathbf{x}_{i}\odot\mathbf{x}_{j}$ (Faskowitz et al., 2020). Reference (Santoro et al., 2023, 2024) extended this approach to characterize $k$-order interactions when $k+1$ nodes simultaneously deviate from their temporal mean activation at time $t$. Although these techniques based on Pearson correlation have provided insights into the brain’s dynamic HOIs, the methods have generally been limited by the lack of adequate temporal resolution (Shine et al., 2015). Now we give a new way to calculate co-fluctuations between several ROIs.
@@ 3.1.3 Construct two signed weighted simplicial complexes
It is important to acknowledge that in a weighted simplicial complex, denoted by $\mathcal{K}^{t}$, there exist two categories of signs for a simplex: concordant signs of positive and discordant signs of negative. The present study focuses exclusively on concordant signs of positive, eschewing the examination of discordant signs of negative. This is due to the potential for such signs to represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). The concordant interactions were subsequently categorised into two distinct groups: those that were positively synergistic and those that were negatively synergistic. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer valuable insights into the complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. As demonstrated in Fig. 2, it exhibits six distinct signed quadruplet HOIs. For each temporal instance $t$, the weighted simplicial complexes are further extracted to construct two signed weighted simplicial complexes $\mathcal{K}^{t}_{positive}$ and $\mathcal{K}^{t}_{negative}$, namely the positive and negative complexes, respectively.
@@ 3.2.2 Homological scaffold and edge projection
At each time point $t$, five distinct weighted networks are constructed: one derived from edge structures representing lower-order interactions, two from signed good tetrahedral structures signifying HOIs, and another two from signed 2D holes illustrating the data’s intrinsic geometric shape from a higher-dimensional perspective. By averaging across the entire time period, the resulting networks exhibit enhanced stability and mitigate the impact of fMRI noise.
@@ 3.3.4 Attention-Guided Feature fusion Mechanism for Classification
For embedding at the graph level ${F}_{i}$ of each channel obtained by flattening ${Z}^{G}_{i}$ ($i=1,2,3$), we ignite an attention-guided feature fusion mechanism to emphasize discriminative features while suppressing redundant information. The formula can be described as:
@@ 4.1.3 Implementation Details
    In HOI-Brain, we employ a Julia-based implementation of the Ripserer and PersistenceDiagrams libraries (Bauer, 2021) to compute persistent homology and homological scaffold, extracting higher-order topological features.
@@ 4.2 Model Comparison
Table,2, 3, 4, 5 demonstrate the comparison results on ADNI, TaoWu, PPMI, and ABIDE datasets. The mean and standard deviation of ten-fold cross-validation were listed, with the best results highlighted in bold and the second-best results in underlining. Compared with traditional ML methods, general-purposed GNN-based models for brain network do not show a significant advantage and attain similar performance in most cases. This finding is consistent with the results reported in existing papers. This is because, under the same input conditions, neural network models typically have a larger number of parameters than conventional machine learning methods when utilizing edge features for better interpretability, making them more susceptible to overfitting. Transformer-based models predominantly outperform GNN-based models across all four datasets in most scenarios, highlighting their superior capability to capture the global structure of brain connectomes. HGNN-based models generally do not demonstrate superiority over other models, with notable performance only observed on the PPMI dataset. This phenomenon can be attributed to two main reasons: first, such models introduce excessive parameters, posing a risk of overfitting; second, their k-hop neighbor-based modeling approach fails to effectively capture the co-fluctuation characteristics among multiple brain regions, thereby inadequately representing HOIs. In PH-based models, such as ATPGCN and Brain-HORs, which introduce higher-order loops or connected component features extracted via persistent homology for brain disease diagnosis, demonstrate significant superiority compared to both graph neural network (GNN)-based models and traditional machine learning models that rely solely on edge features. This highlights the necessity of incorporating higher-order topological features in brain disease diagnosis methods.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNN-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating higher-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network Transformer, can effectively and holistically integrate complementary information from heterogeneous higher-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNN-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of HOIs through the Multiplication of Temporal Derivatives and the extraction of topological features using persistent homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and voids (2D holes) formed by triangular faces and quadruplet.
@@
These experimental results demonstrate the effectiveness of our framework. To the best of our knowledge, our HOI-brain is the first method that distinguishes between positively and negatively synergistic HOIs, and the first attempt to utilize quadruplet-level interaction signatures and two-dimensional void descriptors extracted via Persistent Homology to enhance diagnostic efficacy.
@@ 4.3.1 Ablation study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and voids (2D holes) formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extract lower-order edge features, violating triangles, one-dimensional loops. In addition, we extract unsigned good quadruplets and unsigned two-dimensional voids based our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Table 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be due to the fact that a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true HOIs. Additionally, distinguishing between positively and negatively synergistic HOIs to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may due to the fact that it can offer detailed insights into the communication within the brain.
@@
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network Transformer. The different variants are: not adding signed higher-order features decoupling mechanism (wo-signed), not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Fig. 4, we can see that adding signed higher-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain wins all the cases over the variant (wo-signed) or (wo-fusion). This result is consistent with our assumption that, by introducing the attention mechanism, our model can adaptively fuse positive and negative information as well as information from different channels, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both low-level and high-level feature patterns, which often reside within distinct functional modular structures.
@@ 4.4.1 In-depth Analysis of Attention Mechanism
Next, we also conducted a visualization analysis of the attention scores $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$ derived from the signed higher-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Fig. 7 shows that across both datasets, the importance of negatively synergistic quadruplets is generally greater than that of positively synergistic information in most cases, while the importance of positively synergistic voids exceeds that of negatively synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
In order to conduct a more in-depth investigation into the aforementioned phenomena in the attention mechanism, inter-group comparisons were conducted of the signed higher-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors. These were captured by the model in both the patient and healthy control groups. The ADNI dataset was used as an example. As demonstrated in Fig. 8, which presents a comparison of group differences in higher-order topological features, it was observed that the group differences in the quadruplet-level interaction signatures were found to be highly significant, with all statistical significance levels at $p<0.001$ compared to void descriptors. This finding is consistent with the results of our model visualisation. This phenomenon may be attributed to the disruption of HOIs within the brain during the progression of the disease, which in turn exerts an influence on the high-dimensional hole structures present within the brain. Furthermore, as the disease progresses, the number of positive quadruplets exhibits a gradual decrease, while the number of negative quadruplets initially decreases and subsequently increases. This finding suggests that during the transition from CN to MCI, there is a decline in synergistic interactions in the brain, whether positive or negative. Conversely, as the condition progresses from MCI to AD, brain regions have been observed to exhibit negatively synergistic interactions, concurrently reducing their excitability to preserve the functionality of diverse neural modules. With regard to the number of voids, although no clear trend was observed in the progression of the disease, it was found that positive voids exhibited more significant inter-group differences compared to negative voids. This finding lends further support to the hypothesis that they exhibit higher levels of attention scores. Consequently, CN appears to utilise positive information to construct the whole-brain topological structure, whereas AD relies more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@ 4.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Fig. 9, 10. The attention scores are averaged across all subjects in the ADNI or ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both low-order and higher-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrices on the ADNI or ABIDE dataset in Fig. 11. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and higher-order patterns, which further demonstrates that varying degrees of modularity in both low-order and higher-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.